OpenStudy (llroberts43):

I am stumped on the slope intercept form. 3x-4y = 12. I don't understand why the answer is 3/4x - 3. How does it become x?

Nnesha (nnesha):

you should keep the x slope intercept form is y=mx+b where m is slope and b is y-intercept

Nnesha (nnesha):

you need to solve for y\[\huge\rm 3x-4y=12\] move the `3x`to the right side and then -4

Nnesha (nnesha):

ho would you cancel out the 3x from the left side ?

OpenStudy (llroberts43):

3x - 4y = 12 3x - 3x -4y = -3x +12 -4y and then I don't know what do next.

Nnesha (nnesha):

great! :=)

Nnesha (nnesha):

now how would you cancel out the -4 from the left side ?

OpenStudy (llroberts43):

4y/4?

Nnesha (nnesha):

nice but don't forget it's negative so -4 :=)

OpenStudy (llroberts43):

Yes I saw that but I had already hit post. So I've cancelled out 4 but I still don't understand how the answer becomes 3/4x -3. Where did the 12 go?

Nnesha (nnesha):

\[\huge\rm y=\frac{ 3 x}{ 4 }+\frac{ 12 }{ -4 }\] 12/-4 = ? you can divide it right!

OpenStudy (llroberts43):

Oh....that would equal -3. Right in front of me, thank you!!

Nnesha (nnesha):

:=) my pleasure great job!